OpenStudy (mathmath333):

How many 4 digit numbers are there whose decimal notation contains not more than 2 distinct digits ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ How many 4 digit numbers are there whose decimal notation}\hspace{.33em}\\~\\ & \normalsize \text{ contains not more than 2 distinct digits ? }\hspace{.33em}\\~\\ & a.)\ 672 \hspace{.33em}\\~\\ & b.)\ 576 \hspace{.33em}\\~\\ & c.)\ 360 \hspace{.33em}\\~\\ & d.)\ 448 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (drigobri):

B) 576 is the answer

OpenStudy (mrnood):

how do we know you are right? a straight answer like this is no use to anyone please explain how you derived it, or better still give the poster the method , and let HIM derive it

Parth (parthkohli):

Note to self: No silly mistakes. No silly mistakes. No silly mistakes.

imqwerty (imqwerty):

well what do u mean by decimal notation here :/ sry u posted one more question like decimal notation i got stuck there too

imqwerty (imqwerty):

sry to ask such a silly ques

OpenStudy (mathmath333):

decimal notaion ={0, 1,2,3,4,5,6,7,8,9,} i think

imqwerty (imqwerty):

http://prntscr.com/8kc5b1 :/

Parth (parthkohli):

If there is only one unique digit then there are 9 such numbers (1111, 2222, ...) If there are two unique digits: First, consider all numbers where zero does not appear. (i) Three of one digit and one of the other one. \(9\cdot \binom{4}{3} \cdot 8\) (ii) Two of each digit \(9\cdot \binom{4}{2}\cdot 8\) Now consider those numbers where zero does appear. (i) Three zeroes and one other digit. This can only be done in 9 ways (1000, 2000, ...) (ii) Two zeroes and two of the other digit. This can be done in \(9\cdot \binom{3}{2}\)

Parth (parthkohli):

I think "decimal notation" means nothing but base 10 integers here.

Parth (parthkohli):

(iii) One zero and three of other digits.\[3\cdot 9~ways \]

Parth (parthkohli):

wow why am I getting 765

Parth (parthkohli):

@ganeshie8 !!!

Parth (parthkohli):

Can you spot any mistake?

OpenStudy (mathmath333):

i think u double counted something

Parth (parthkohli):

What?

Parth (parthkohli):

I'm actually getting 765 + 27 = 792 :P

ganeshie8 (ganeshie8):

Slightly modified : If there is only one unique digit then there are 9 such numbers (1111, 2222, ...) If there are two unique digits: First, consider all numbers where zero does not appear. (i) Three of one digit and one of the other one. \(9\cdot \binom{4}{3} \cdot 8\) (ii) Two of each digit \(9\cdot \dfrac{\binom{4}{2}}{\color{red}{2}}\cdot 8\) Now consider those numbers where zero does appear. (i) Three zeroes and one other digit. This can only be done in 9 ways (1000, 2000, ...) (ii) Two zeroes and two of the other digit. This can be done in \(9\cdot \binom{3}{2}\) (iii) One zero and three of the other digit. This can be done in \(\color{red}{9\cdot \binom{3}{1}}\) @ParthKohli

Parth (parthkohli):

why would you change \(\binom{3}2\) to \(\binom{3}1\) and good catch